Charlie Corp went “public” via an IPO that was completed earlier in the day. The IPO resulted in 40%

of the company’s 5,000,000 total shares being sold to institutional investors, which in turn made the

shares available for purchase on the NYSE. The IPO price was $40 per share. The stock began

trading at $46 per share, and closed at $44 per share.

a. How much capital did the company raise?

b. How much did institutional investors make when the stock opened?

c. What was Charlie Corp’s total market value when the stock closed?
Answer #1
  1. ) Amount of capital= 40 % x 5000000 x $ 40

= ( 40 / 100 ) x 5000000 x $ 40 = $ 80000000 =$ 80 million

b) The amount that institutional investors make when the stock opened = (40/100) x 5000000 x ($ 46 - $ 40)
= (40 / 100) x 5000000 x $ 6
= $ 12 million
This is the amount the institutional investors will make if they immediately sell the shares when they open at $ 46 /share.

If the institutional investors does not sell the total amount when it opens = (40/100) x 5000000 x $ 46
= $ 92000000
= $ 92 million

c) The stock closed at $ 44 / share.
Charlie Corp's total market value when the stock closed = 5000000 x $ 44
= $ 220 million